# Generalized Pinching-Antenna Systems: A Tutorial on Principles, Design Strategies, and Future Directions

### v1

@@ Abstract
Pinching-antenna systems have emerged as a novel and transformative flexible-antenna architecture for next-generation wireless networks. They offer unprecedented flexibility and spatial reconfigurability by enabling dynamic positioning and activation of radiating elements along a signal-guiding medium (e.g., dielectric waveguides), which is not possible with conventional fixed antenna systems. In this paper, we introduce the concept of generalized pinching antenna systems, which retain the core principle of creating localized radiation points on demand, but can be physically realized in a variety of settings. These include implementations based on dielectric waveguides, leaky coaxial cables, surface-wave guiding structures, and other types of media, employing different feeding methods and activation mechanisms (e.g., mechanical, electronic, or hybrid). Despite differences in their physical realizations, they all share the same inherent ability to form, reposition, or deactivate radiation sites as needed, enabling user-centric and dynamic coverage. We first describe the underlying physical mechanisms of representative generalized pinching-antenna realizations and their associated wireless channel models, highlighting their unique propagation and reconfigurability characteristics compared with conventional antennas. Then, using dielectric waveguide-based pinching antennas as primary examples, we review several representative pinching-antenna system architectures, ranging from single- to multiple-waveguide configurations, and discuss advanced design strategies tailored to these flexible deployments. Furthermore, we examine their integration with emerging wireless technologies to enable synergistic, user-centric solutions. Finally, we identify key open research challenges and outline future directions, charting a pathway toward the practical deployment of generalized pinching antennas in next-generation wireless networks.
@@ I-A Background and Motivations
The evolution of wireless networks is entering a new era with the advent of sixth-generation (6G) wireless systems. Exceeding the capabilities of previous generations, 6G is expected to support various new applications and use cases, including immersive extended reality, holographic telepresence, intelligent robotics, edge artificial intelligence (AI), generative AI, and large-scale Internet of Things (IoT) deployments [1, 2, 3]. These new use cases have different requirements on data rates, latency, reliability, and connectivity density. Meeting these diverse requirements requires innovations across the protocol stack. In the physical layer, one of the most critical components is the multi-antenna subsystem, which directly influences the system throughput, spatial coverage, and energy efficiency [4, 5]. The wireless industry has traditionally relied on fixed or semi-static multi-antenna systems, such as conventional multiple-input multiple-output (MIMO) arrays. These technologies exploit spatial diversity, enhance spectral efficiency, and combat multipath fading, thereby offering significant performance improvements in the fourth-generation (4G) and fifth-generation (5G) networks. However, a major limitation of traditional multi-antenna solutions is their limited reconfigurability. Once the antennas are fabricated and installed, their physical and electromagnetic properties are largely fixed. Consequently, these systems lack the flexibility to adapt to changes in propagation conditions, spectrum usage, or user locations. Looking ahead, future wireless networks will become increasingly dense and dynamic. They will need to cope with users’ mobility, reconfigurable environments, and highly variable spectrum allocations. The limited reconfigurability of traditional antennas may lead to significant performance bottlenecks in such scenarios.
@@
This context has catalyzed a major shift toward flexible antenna architectures. Representative examples include reconfigurable intelligent surfaces (RISs), also known as intelligent reconfigurable surfaces (IRSs) [6, 7], fluid antennas [8, 9], and movable antenna systems [10, 11]. These systems have attracted significant research attention due to their ability to dynamically modify wireless channels and enhance communication performance. RISs electronically adjust the reflection coefficients of large arrays of passive elements to enable programmable propagation environments. Fluid antenna systems use the physical properties of conducting fluids to alter the antenna’s position or geometry, enabling agile adaptation to changing signal conditions. Movable antenna systems provide flexibility by mechanically displacing antenna elements within a predefined region to support position-dependent channel optimization. While these techniques bring significant advances, they cannot directly address the challenging issue of line-of-sight (LoS) blockage. Each of the aforementioned techniques has its own potential practical limitations. RIS-based systems may suffer from double attenuation in the transmitter-to-RIS and RIS-to-receiver paths. This can result in substantial path loss, especially under non-line-of-sight (NLoS) conditions. Both fluid and movable antennas generally restrict their movement to a few wavelengths, limiting their ability to create new LoS channels and mitigate large-scale path loss when user locations vary significantly. These potential limitations affect their performance in highly dynamic or densely deployed network scenarios.
@@
To address the aforementioned challenges, the research community is exploring advanced reconfigurable antenna technologies beyond current paradigms. Among the emerging technologies, pinching-antenna systems have recently received significant attention as a promising new approach. They have the potential of enhancing the flexibility, reducing the implementation cost, providing fine-grained control, and improving integration into next-generation wireless networks.
@@
Figure 1: Conceptual illustration of generalized pinching-antenna systems: (a) Dielectric waveguide-based pinching antennas use discrete dielectric particles to enable localized radiation along a dielectric waveguide; (b) LCX-based pinching antennas employ periodic slots on coaxial cables, with electronic switches enabling on-demand signal leakage through selective segment activation; and (c) Pinching-inspired antenna, in which active antennas are strategically placed along a transmission surface to enable user-centric and on-demand wireless coverage.
@@ I-B Generalized Pinching-Antenna Systems: Basic Principles and Key Advantages
Generalized pinching-antenna systems represent a broad and evolving class of flexible-antenna technologies that enable dynamic and on-demand creation or reconfiguration of radiating points along a transmission medium. The fundamental principle is to enable localized electromagnetic energy leakage at specific locations, which allows wireless coverage to be tailored according to user demand or scenario requirements. This approach overcomes the limitations of traditional fixed antenna installations by offering better adaptability and user-centric deployment in wireless networks. In Fig. 1, three representative examples of generalized pinching-antenna systems are presented.
@@
The representative example of generalized pinching antennas is NTT DOCOMO’s dielectric waveguide-based pinching antenna [12], where the radiating elements are physically pinched (i.e., attached) to a dielectric waveguide using discrete dielectric particles or clips (see Fig. 1(a)). Each pinch disrupts signals, which are propagated along the waveguide, causing controlled energy leakage at the desired position and effectively “creating” an antenna. This process is reversible and reconfigurable. That is, the radiating elements can be deployed, repositioned, or removed flexibly, enabling highly adaptive coverage with minimal infrastructure modification [13].
@@
Leaky coaxial cables (LCXs) are another important example of generalized pinching antennas. Conventional LCX antennas radiate electromagnetic signals through periodic slots in coaxial cables, providing reliable wireless coverage in tunnels, subway systems, mines, and other confined or elongated environments [14, 15, 16]. Compared to DOCOMO’s pinching antennas, conventional LCX antennas lack flexibility and user-centric control, as their radiation points are fixed and cannot be dynamically reconfigured to satisfy the changing coverage requirement. However, by dividing an LCX cable into multiple segments and allowing these segments to be selectively activated or deactivated, e.g., through electronic switching, radiating sections can be dynamically tailored to user locations or coverage needs, which yields LCX-based pinching antennas as shown in Fig. 1(b). In particular, this localized and user-centric radiation mechanism parallels the on-demand flexibility of dielectric waveguide-based pinching antennas, albeit through a different physical implementation.
@@
The key idea of pinching antennas can also be used to design various pinching-inspired antenna arrays. Another representative example is shown in Fig. 1(c), where the signals are first guided through a metal pipe via the surface-wave mode and then radiated in an on-demand mode by applying an active antenna close to the served user [17, 18]. Such systems can operate in a manner similar to power track systems, where each pinching antenna functions like a plug-in module that can be flexibly deployed along the medium. Unlike surface-wave communication, which is non-radiant, the use of such pinching-inspired antenna arrays ensures that electromagnetic signals are radiated into free space at the chosen locations. Another example draws from the concept of radio stripes [19, 20], in which a series of radiating sites are embedded along the extended conductive structures. In this approach, signals are transmitted with low attenuation through the radio stripe to positions that are close to the intended users, where they are subsequently radiated into free space. Integrating active antennas, directional couplers, or segmented elements enables these antennas to activate radiation on demand and adapt dynamically to coverage and user demands.
@@
  • Practicality, Cost Effectiveness, and Simplicity: Generalized pinching-antenna systems offer a highly practical and economically viable solution by leveraging low-cost transmission medium and readily manufacturable pinching units. Unlike conventional antenna arrays that require complex radio frequency (RF) chains for each radiating element, pinching antennas significantly reduce hardware complexity by enabling signal radiation directly from waveguide pinches. This minimalist architecture not only reduces material and deployment costs but also simplifies system integration, maintenance, and scalability, making it particularly attractive for large-scale, user-centric, and flexible wireless deployments in both indoor and outdoor scenarios.
An illustration of the advantages of generalized pinching antennas is provided in Fig. 2.
TABLE I: Outline of the paper.
@@ C. Pinching-Inspired Antenna Systems
C. Uplink Pinching-Antenna Systems
@@ C. Synergizing Pinching Antennas with Other Advanced Techniques
E. Pinching-Antenna-Enabled Extremely Large-Scale Antenna Array F. Pinching Antenna-Assisted Wireless Edge Intelligence
G. Robust Pinching-Antenna System Design H. Real-Time Low-Complexity Control of Mechanical Antenna Placement
@@ I-C Contributions and Organization of the Paper
There has been a surge of research interest in pinching-antenna systems recently. Existing studies, including several magazine and tutorial papers [21, 22, 23] and the references therein, have significantly advanced the understanding and design of such systems. Nevertheless, these works have primarily concentrated on dielectric waveguide–based implementations and have not offered a unified or systematic treatment that extends the pinching concept to other possible physical realizations.
In contrast, this paper is the first to formally introduce the concept of generalized pinching-antenna systems. This extended framework preserves the core idea of creating localized radiation points on demand but encompasses a variety of physical implementations, including dielectric waveguides, LCXs, and surface-wave guiding structures, together with diverse feeding methods and activation mechanisms. We provide a comprehensive tutorial that covers their fundamental principles, representative architectures, design strategies, integration with emerging wireless technologies, and open research challenges. The main contributions of this paper are as follows:
@@
  • We then focus on DOCOMO’s pinching antennas and review several representative system architectures, including single- and multi-waveguide configurations. We further survey advanced design strategies for radio resource allocation, beamforming, and antenna placement tailored to pinching-antenna deployments. In addition, we highlight specific system settings where the design problem admits closed-form analytical solutions, which not only enable low-complexity designs but also provide fundamental insights into pinching-antenna systems.
@@
  • We outline several promising research directions and open challenges for future research on pinching-antenna systems. These include pinching-antenna system design under LoS blockage, the realization and optimization of efficient and controllable pinching-antenna radiators, and the development of accurate and credible electromagnetic models, the exploitation of near-field communication, system optimization under user position uncertainty, and scalable and real-time control strategies.
TABLE II: List of acronyms.
@@
This paper is organized as follows. Section II introduces the fundamental principles of generalized pinching-antenna systems, including their physical realizations, signal propagation mechanisms, and unique channel characteristics. Representative system architectures and design methodologies for pinching-antenna deployments are systematically surveyed in Section III, where both single- and multi-waveguide systems and corresponding optimization strategies are covered. Section IV explores the different applications of pinching antennas, detailing their roles in ISAC, security, wireless power transfer, edge computing, and other advanced wireless paradigms. Section V outlines several open questions and research directions. Finally, conclusions are drawn in Section VI. Table I presents the outline of this paper and Table II lists the acronyms used.
Notations: Column vectors and matrices are denoted by boldfaced lowercase and uppercase letters, e.g., $\bf x$} and $\bf X$}, respectively. $\mathbb{C}^{N}$ denotes the set of $N$-dimensional complex vectors. Superscripts $(\cdot)^{\top}$ and $(\cdot)^{\mathsf{H}}$ denote the transpose and Hermitian operations, respectively. ${\bf x}||^{2}$ denotes the square of the Euclidean norm of vector $\bf x$}}. $\exp(\cdot)$ denotes the exponential function. ${\mathbb{E}}_{n}[\cdot]$ represents the statistical expectation operation with respect to the random variable $n$.

### v2

@@ Abstract
Pinching-antenna systems have emerged as a novel and transformative flexible-antenna architecture for next-generation wireless networks. They offer unprecedented flexibility and spatial reconfigurability by enabling dynamic positioning and activation of radiating elements along a signal-guiding medium (e.g., dielectric waveguide), which is not possible with conventional fixed antenna systems. In this paper, we introduce the concept of generalized pinching antenna systems, which retain the core principle of creating localized radiation points on demand, but can be physically realized with a variety of technologies. These include implementations based on dielectric waveguides, leaky coaxial cables, surface-wave guiding structures, and other types of media, employing different feeding methods and activation mechanisms (e.g., mechanical, electronic, or hybrid). Despite differences in their physical realizations, they all share the same inherent ability to form, reposition, or deactivate radiation sites as needed, enabling user-centric and dynamic coverage. We first describe the underlying physical mechanisms of representative generalized pinching-antenna realizations and their associated wireless channel models, highlighting their unique propagation and reconfigurability characteristics compared with conventional antennas. Then, using dielectric waveguide-based pinching antennas as primary examples, we review several representative pinching-antenna system architectures, ranging from single- to multiple-waveguide configurations, and discuss advanced design strategies tailored to these flexible deployments. Furthermore, we examine their integration with emerging wireless technologies to enable synergistic, user-centric solutions. Finally, we identify key open research challenges and outline future directions, charting a pathway toward the deployment of generalized pinching antennas in next-generation wireless networks.
@@ I-A Background and Motivation
The evolution of wireless networks is entering a new era with the advent of sixth-generation (6G) wireless systems. Exceeding the capabilities of previous generations, 6G is expected to support various new applications and use cases, including immersive extended reality, holographic telepresence, intelligent robotics, edge artificial intelligence (AI), generative AI, and large-scale Internet of Things (IoT) deployments [1, 2, 3, 4, 5, 6]. These new use cases have different requirements on data rates, latency, reliability, and connectivity density. Meeting these diverse requirements requires innovations across the protocol stack. In the physical layer, one of the most critical components is the multi-antenna subsystem, which directly influences system throughput, spatial coverage, and energy efficiency [7, 8]. The wireless industry has traditionally relied on fixed or semi-static multi-antenna systems, such as conventional multiple-input multiple-output (MIMO) arrays. These technologies exploit spatial diversity, enhance spectral efficiency, and combat multipath fading, thereby enabling significant performance improvements in fourth-generation (4G) and fifth-generation (5G) wireless networks. However, a major limitation of traditional multi-antenna solutions is their limited reconfigurability. Once the antennas are fabricated and installed, their physical and electromagnetic properties are largely fixed. Consequently, these systems lack the flexibility to adapt to changes in propagation conditions, spectrum usage, and user locations. Looking ahead, future wireless networks will become increasingly dense and dynamic. They will need to cope with users’ mobility, reconfigurable environments, and highly variable spectrum allocations. The limited reconfigurability of traditional antennas may lead to significant performance bottlenecks in such scenarios.
@@
These challenges have catalyzed a major shift toward flexible antenna architectures. Representative examples include reconfigurable intelligent surfaces (RISs), also known as intelligent reconfigurable surfaces (IRSs) [9, 10], fluid antennas [11, 12], and movable antenna systems [13, 14]. These systems have attracted significant research attention due to their ability to dynamically modify wireless channels and enhance communication performance. RISs electronically adjust the reflection coefficients of large arrays of passive elements to enable programmable propagation environments. Fluid antenna systems use the physical properties of conducting fluids to alter the antennas’ positions or geometry, enabling agile adaptation to changing signal conditions. Movable antenna systems provide flexibility by mechanically displacing antenna elements within a predefined region to support position-dependent channel optimization. While these techniques bring significant advances, they cannot directly address the challenging issue of line-of-sight (LoS) blockage. Each of the aforementioned techniques has its own practical limitations. RIS-based systems may suffer from double attenuation in the transmitter-to-RIS and RIS-to-receiver paths. This can result in substantial path loss, especially under non-line-of-sight (NLoS) conditions. Both fluid and movable antennas generally restrict their movement to a few wavelengths, limiting their ability to create new LoS channels and mitigate large-scale path loss when user locations vary significantly. These limitations may affect their performance in highly dynamic or densely deployed network scenarios.
@@
To address the aforementioned challenges, the research community is exploring advanced reconfigurable antenna technologies beyond current paradigms. Among the emerging technologies, pinching-antenna systems have recently received significant attention as a promising new approach. They have the potential to enhance flexibility, reduce implementation cost, provide fine-grained control, and improve integration into next-generation wireless networks.
@@
Figure 1: Conceptual illustration of generalized pinching-antenna systems: (a) Dielectric waveguide-based pinching antennas use discrete dielectric particles or clips to enable localized radiation along a dielectric waveguide; (b) LCX-based pinching antennas employ periodic slots on coaxial cables, with electronic switches enabling on-demand signal leakage through selective segment activation; and (c) Pinching-inspired antennas, in which active antennas are strategically placed along a transmission surface to support user-centric and on-demand wireless coverage.
@@ I-B Generalized Pinching-Antenna Systems: Basic Principles and Key Advantages
Generalized pinching-antenna systems represent a class of antenna architectures in which radio frequency (RF) signals are first guided along a physical medium and then are selectively radiated into free space at configurable radiation points, creating wireless links near the users. The fundamental principle is to enable localized electromagnetic energy leakage at specific locations, which allows wireless coverage to be tailored according to user demand and scenario requirements. This approach overcomes the limitations of traditional fixed antenna installations by offering better adaptability and user-centric deployment in wireless networks. In Fig. 1, three representative examples of generalized pinching-antenna systems are presented. This includes NTT DOCOMO’s dielectric waveguide-based pinching-antenna system [15], where the radiating elements are physically pinched (i.e., attached) to a dielectric waveguide using discrete dielectric particles or clips (see Fig. 1(a)). Each pinch disrupts signals, which are propagated along the waveguide, causing controlled energy leakage at the desired position, effectively “creating” an antenna. This process is reversible and reconfigurable. That is, the radiating elements can be deployed, repositioned, or removed flexibly, enabling highly adaptive coverage with minimal infrastructure modification [16].
@@
Leaky coaxial cables (LCXs) are another important example of generalized pinching antennas. Conventional LCX antennas radiate electromagnetic signals through periodic slots in coaxial cables, providing reliable wireless coverage in tunnels, subway systems, mines, and other confined or elongated environments [17, 18, 19, 20]. Compared to DOCOMO’s pinching antennas, conventional LCX antennas lack flexibility and user-centric control, since their radiation points are fixed and cannot be dynamically reconfigured to satisfy the changing coverage requirement. However, by dividing an LCX cable into multiple segments and allowing these segments to be selectively activated or deactivated, e.g., through electronic switching, radiating sections can be dynamically tailored to user locations or coverage requirements, which yields an LCX-based pinching-antenna system as shown in Fig. 1(b). This localized and user-centric radiation mechanism parallels the on-demand flexibility of dielectric waveguide-based pinching antennas, albeit through a different physical implementation.
@@
The key idea of pinching antennas can also be used to design various pinching-inspired antenna arrays. Another representative example is shown in Fig. 1(c), where the signals are first guided through a metal pipe via the surface-wave mode and then radiated in an on-demand manner by applying an active antenna close to the users [21, 22]. Such systems can operate similarly to power track systems, where each pinching antenna functions like a plug-in module that can be flexibly deployed along the medium. Unlike surface-wave communication, which is non-radiating, the use of such pinching-inspired antenna arrays ensures that electromagnetic signals are radiated into free space at the chosen locations. Another example draws from the concept of radio stripes [23, 24], in which a series of radiating sites are embedded along the extended conductive structures. In this approach, signals are transmitted with low attenuation through the radio stripe to positions that are close to the intended users, where they are subsequently radiated into free space. Integrating active antennas, directional couplers, or segmented elements enables these antennas to activate radiation on demand and adapt dynamically to coverage and user demands.
@@
  • Practicality, Cost Effectiveness, and Simplicity: Generalized pinching-antenna systems offer a highly practical and economically viable solution by leveraging low-cost transmission media and readily manufacturable pinching units. Unlike conventional antenna arrays that require complex RF chains for each radiating element, pinching antennas significantly reduce hardware complexity by enabling signal radiation directly from waveguide pinches. This minimalist architecture not only reduces material and deployment costs but also simplifies system integration, maintenance, and scalability, making it particularly attractive for large-scale, user-centric, and flexible wireless deployments in both indoor and outdoor scenarios.
TABLE I: Qualitative comparison of pinching-antenna systems, RIS, and fluid/movable antennas.
Aspect Generalized pinching antennas RIS Fluid/movable antennas
Signal propagation Wired segment ++ short-distance wireless link Two-hop BS–RIS–user wireless link Single-hop wireless link with reconfigurable radiating point
Path-loss behaviour Wired segment exhibits low loss along guided structures, while short wireless link incurs only small free-space path loss Experiences two-hop path loss; performance sensitive to BS–RIS–user topology Similar to conventional antennas, path loss is mainly determined by user–BS link distance, with gains from local fading diversity
Reconfiguration scale Large-scale along guiding media (tens of metres, many radiating points) Panel-scale on a fixed surface (element-wise reflection control) Small-scale at wavelength-level antenna reconfiguration and movement
Hardware / deployment Dielectric waveguides, LCX, or similar media; few RF chains feeding many radiating points Large passive panel on existing surfaces; no extra RF chains but many tunable elements Fluid antennas use tunable conductive liquids in small cavities, while movable antennas rely on compact mechanical actuators.
TABLE II: Comparison between this work and representative references on pinching-antenna systems.
Reference Paper Type Pinching-antenna realizations Main elements covered
Dielectric waveguide- based LCX- based Pinching- inspired Channel modelling and system optimization Impact of key system parameters (e.g., attenuation, blockage) Integration with advanced techniques, applications, and open directions
[16] Technical
[25] Magazine
[26] Magazine
[27] Tutorial
This work Tutorial
An illustration of the advantages of generalized pinching antennas is provided in Fig. 2. As shown from the figure, generalized pinching-antenna systems operate in a different regime compared to both RISs and fluid/movable antennas. RISs reshape the small-scale wireless channel using a fixed reflecting surface by tuning a large number of passive elements, while fluid/movable antennas reconfigure a radiating point within a small region to exploit local fading diversity. In both cases, the large-scale propagation geometry and the base station (BS)–user path loss remain essentially unchanged. In contrast, generalized pinching antennas exploit low-loss guiding media (dielectric waveguides, LCX, or surface-wave conductors) to realize large-scale channel reconfiguration employing a small number of feeder points, creating, removing, or repositioning radiating spots over tens of meters along installed structures, while incurring only modest RF-chain and cabling overhead. A qualitative comparison between these flexible-antenna techniques is summarized in Table I.
TABLE III: Outline of the paper.
@@ C. Pinching-Inspired Antenna Systems
C. Pinching-Antenna System Design under Random LoS and NLoS Channels D. Uplink Pinching-Antenna Systems
E. Control and Implementation of Pinching-Antenna Positioning
@@ C. Synergizing Pinching Antennas with Other Advanced Techniques
E. Environment-Aware Pinching-Antenna System Designs F. Pinching-Antenna-Enabled Extremely Large-Scale Antenna Array
G. Pinching Antenna-Assisted Wireless Edge Intelligence H. Robust Pinching-Antenna System Design
I. Real-Time Low-Complexity Control of Mechanical Antenna Placement J. Practical Implementation Challenges and Scalability
@@ I-C Contributions and Organization of the Paper
There has been a surge of research interest in pinching-antenna systems recently. Existing studies, including several technical, magazine, and tutorial papers on dielectric waveguide-based implementations, see [16, 25, 26, 27] and the references therein, have significantly advanced the understanding and design of such systems. However, as summarized in Table II, these works mainly focus on a single physical realization, i.e., dielectric waveguides, and address selected aspects in isolation, without offering a unified or systematic treatment that extends the pinching-antenna concept to other realizations or systematically examines the impact of key system parameters on system performance.
This paper is the first to formally introduce the concept of generalized pinching-antenna systems. This extended framework preserves the core idea of creating localized radiating points on demand but encompasses a variety of physical implementations, including dielectric waveguides, LCXs, and surface-wave guiding structures, together with diverse feeding methods and activation mechanisms. We provide a comprehensive tutorial that covers their fundamental principles, representative architectures, design strategies, integration with emerging wireless technologies, and open research challenges. The main contributions of this paper are as follows:
@@
  • Using DOCOMO’s pinching antennas as an example, we review several representative system architectures, including single- and multi-waveguide configurations. We further survey advanced design strategies for radio resource allocation, beamforming, and antenna placement tailored to pinching-antenna deployments. In addition, we highlight specific system settings where the design problem admits closed-form analytical solutions, which not only enable low-complexity designs but also provide fundamental insights into pinching-antenna systems.
@@
  • We outline several promising research directions and open challenges for future research on pinching-antenna systems. These include pinching-antenna system design under LoS blockage, the realization and optimization of efficient and controllable pinching-antenna radiating points, and the development of accurate and credible electromagnetic models, the exploitation of near-field communication, system optimization under user position uncertainty, and scalable and real-time control strategies.
TABLE IV: List of acronyms.
@@
This paper is organized as follows. Section II introduces the fundamental principles of generalized pinching-antenna systems, including their physical realizations, signal propagation mechanisms, and unique channel characteristics. Representative system architectures and design methodologies for pinching-antenna deployments are systematically surveyed in Section III, where both single- and multi-waveguide systems and the corresponding optimization strategies are covered. Section IV explores different applications of pinching antennas, detailing their roles in ISAC, security, wireless power transfer, edge computing, and other advanced wireless paradigms. Section V outlines several open questions and research directions. Finally, conclusions are drawn in Section VI. Table III presents the outline of this paper and Table IV lists the acronyms used.
Notations: Column vectors and matrices are denoted by boldfaced lowercase and uppercase letters, e.g., $\bf x$} and $\bf X$}, respectively. $\mathbb{C}^{N}$ denotes the set of $N$-dimensional complex vectors. Superscripts $(\cdot)^{\top}$ and $(\cdot)^{\mathsf{H}}$ denote the transpose and Hermitian operations, respectively. ${\bf x}||^{2}$ denotes the square of the Euclidean norm of vector $\bf x$}}. $\exp(\cdot)$ denotes the exponential function. $\mathbb{P}(\cdot)$ denotes the probability of an event. ${\mathbb{E}}_{n}[\cdot]$ represents the statistical expectation operation with respect to the random variable $n$.